OpenStudy (jessiegonzales):

MEDAL! 14. Solve the absolute value equation. l2x-1l=3 A. x=2 or x=-1 B. x=1 or x=2 C. x-1 or x=-1 D. x=1 or x=4

OpenStudy (welshfella):

solve the 2 equations 2x - 1 = 3 and 2x - 1 = -3

OpenStudy (jessiegonzales):

whats the first step?

OpenStudy (welshfella):

what s the first step to solve 2x -1 = 3?

OpenStudy (jessiegonzales):

divide

OpenStudy (welshfella):

you need to first isolate 2x so you add 1 to both sides of the equation

OpenStudy (welshfella):

then you divide both sides by 2 and you do exactly the same process with the second equation

OpenStudy (jessiegonzales):

3x-1=4

OpenStudy (welshfella):

hope this helps gotta go right now

OpenStudy (jessiegonzales):

does that look right

OpenStudy (welshfella):

2x - 1 = 3 +1 +1

OpenStudy (jessiegonzales):

2x=4

OpenStudy (welshfella):

gotta go

OpenStudy (jessiegonzales):

x=2

OpenStudy (jessiegonzales):

answer is A

OpenStudy (welshfella):

yes